# Bridging Entanglement and Magic Resources through Operator Space

### v1

@@ Appendix B Average-case lower bound on LOE from unitary ensembles
i.e. the purity of the Choi stat $|O_{U}\rangle\!\rangle$. Above, we consider operator entanglement to be across any (spatial) bipartition of $\mathcal{H}_{A}\otimes\mathcal{H}_{\bar{A}}$, of $N_{A}$ and $N_{\bar{A}}$ qubits respectively. In the case of both unitary ensembles considered here, the particular qubits in $\mathcal{H}_{A}$ does not matter, only the size $N_{A}$. This is because both the Clifford and unitary groups are invariant under left or right multiplication by a permutation matrix (as permutations are contained in the Clifford group is contained in the unitary group). We can write the operator purity in replica space as
@@
from the bound $E^{(2)}\leq 2E^{(\infty)}$. This final bound can also be used to provide a lower bound in Thm. 2 in terms of Rényi entropies $\alpha>2$, but we do not bother as $S^{(\alpha)}$ for $0\leq\alpha\leq 2$ are more informative Verstraete and Cirac (2006); Schuch et al. (2008).
B.1 Clifford and Haar averages
@@
B.2 $\nu$-compressible ensemble
We first consider the $\nu$-compressible ensemble, as it is somewhat simpler to compute than the $T$-doped averages. Recall that the $\nu$-compressible ensemble is defined as the class of unitaries $\{U_{\nu}=C_{0}(V\otimes\mathds{1}_{N-\nu})C_{1}:C_{0/1}\in\mathcal{C}_{N},V\in\mathcal{U}_{\nu}\}$, equipped with the uniform measure over $C_{0},C_{1}\in\mathcal{C}_{N}$ and $V\in\mathcal{U}_{\lceil\nu/2\rceil}$ on the Clifford and unitary groups respectively. For ease of notation, we will write $V\equiv(V\otimes\mathds{1}_{N-\lceil\nu/2\rceil})$ in the following, and write ${\tilde{\nu}}:=\lceil\nu/2\rceil$. As usual, we take initial operator to be a non-identity Pauli string, $O\in\mathcal{P}_{N}\backslash\{\mathds{1}\}$.
@@ B.2 $\nu$-compressible ensemble
This was computed symbolically in Mathematica. The above is a generalization of Eq. (6) presented in the main text; it is valid for any $D$, $\nu$, and $D_{A}$. We will now explore various limits of this quantity to extract some physical insight.
@@
Note that the final expression agrees with the leading order term for the Haar averaged operator purity from Ref. Kudler-Flam et al. (2021) (see Eq. (29)). Moreover, choosing an equal-sized bipartition of $D_{A}=\sqrt{D}$, we arrive at Eq. (7) in the main text.
@@
Here, we have neglected factors proportional to $1/D^{2}$. To obtain a lower bound as that which appears in Thm. 2, we apply Jensen’s inequality,
@@
where the final line is an approximation for small $\frac{D_{\nu}^{2}}{D}+\frac{1}{D_{\nu}^{2}}$. The exponential of the negative of the LHS is exactly the result of Eq. (6).
B.3 $T$-doped Clifford ensemble
@@
Now, for the matrix $F$, we assume that $K$ is a single qubit $R^{Z}_{\theta}$ gate wlog acting on the first qubit, which we call $P_{\theta}$. For the rest of the proof, we will use the Lemmas 8-9.
@@
Expanding the above expression to leading order in large $D$, and taking the $-\log$, we arrive at Eq. (3),
@@
which is valid for $1/D\ll 1$.

### v2

@@ Appendix B Bounding Rényi LOE from average operator purity in replica space
i.e. the purity of the Choi stat $|O_{U}\rangle\!\rangle$. Above, we consider operator entanglement to be across any (spatial) bipartition of $\mathcal{H}_{A}\otimes\mathcal{H}_{\bar{A}}$, of $N_{A}$ and $N_{\bar{A}}$ qubits respectively. In the case of both unitary ensembles considered here, the particular qubits in $\mathcal{H}_{A}$ does not matter, only the size $N_{A}$. This is because both the Clifford and unitary groups are invariant under left or right multiplication by a permutation matrix (as permutations are contained in the Clifford group, and the Clifford group is contained in the unitary group). We can write the operator purity in replica space as
@@
from the bound $E^{(2)}\leq 2E^{(\infty)}$. This final bound can also be used to provide a lower bound in Thm. 1 in terms of Rényi entropies $\alpha>2$, but we do not bother as $S^{(\alpha)}$ for $0\leq\alpha\leq 2$ are more informative Verstraete and Cirac (2006); Schuch et al. (2008).
Appendix C Clifford and Haar averages
@@
Appendix D Average operator purity over the $\nu$-compressible ensemble
We first consider the $\nu$-compressible ensemble, as it is somewhat simpler to compute than the $T$-doped averages. Recall that the $\nu$-compressible ensemble is defined as the class of unitaries $\{U_{\nu}=C_{0}(V\otimes\mathds{1}_{N-\nu})C_{1}:C_{0/1}\in\mathcal{C}_{N},V\in\mathcal{U}_{\nu}\}$, equipped with the uniform measure over $C_{0},C_{1}\in\mathcal{C}_{N}$ and $V\in\mathcal{U}_{\lceil\nu/2\rceil}$ on the Clifford and unitary groups respectively. For ease of notation, we will write $V\equiv(V\otimes\mathds{1}_{N-\lceil\nu/2\rceil})$ and ${\tilde{\nu}}:=\lceil\nu/2\rceil$ in the following. As usual, we take initial operator to be some non-identity Pauli string, $O\in\mathcal{P}_{N}\backslash\{\mathds{1}\}$.
@@ Appendix D Average operator purity over the $\nu$-compressible ensemble
This was computed symbolically in Mathematica. The above is a generalization of the top upper bound presented in Thm. 1; it is valid for any $D$, $\nu$, and $D_{A}$. We will now explore various limits of this quantity to extract some physical insight.
@@
Note that the final expression agrees with the leading order term for the Haar averaged operator purity from Ref. Kudler-Flam et al. (2021) (see Eq. (29)). Moreover, choosing an equal-sized bipartition of $D_{A}=\sqrt{D}$, we arrive at Eq. (6) in the main text,
@@
Here, we have neglected factors proportional to $1/D^{2}$. To obtain a lower bound as that which appears in Thm. 1, we apply Jensen’s inequality,
@@
where the final line is an approximation for small $\frac{D_{\nu}^{2}}{D}+\frac{1}{D_{\nu}^{2}}$. This is our final bound in terms of unitary nullity.
Appendix E Average operator purity over the $T$-doped Clifford ensemble
@@
Now, for the matrix $F$, we assume that $K$ is a single qubit $R^{Z}_{\theta}$ gate wlog acting on the first qubit, which we call $P_{\theta}$. For the rest of the proof, we will use the Lemmas 7-8.
@@
Expanding the above expression to leading order in large $D$, and taking the $-\log$, we arrive at,
@@
which is valid for $1/D\ll 1$. Applying Jensen’s inequality and the hierarchy of Rényi entropies (see App. B), we arrive at the second lower-bound of Thm. 1.